Listed below are the amounts of net worth (in millions of dollars) of the ten wealthiest celebrities in a country. Construct a 90% confidence interval. What does the result tell us about the population of all celebrities? Do the data appear to be from a normally distributed population as required? 

268 211 186 171 169 156 147 147 147 142 

What is the confidence interval estimate of the population mean μ? 


 What does the result tell us about the population of all celebrities? Select the correct choice below and, if necessary, fill in the answer box(es) to complete your choice. 

A. Because the ten wealthiest celebrities are not a representative sample, this doesn't provide any information about the population of all celebrities. 

B. We are 90% confident that the interval from S million to $ million actually contains the true mean net worth of all celebrities. (Round to one decimal place as needed.) 

C. We are confident that 90% of all celebrities have a net worth between $ million and $ million (Round to one decimal place as needed.) 


Do the data appear to be from a normally distributed population as required? 

A. No, but the points in the normal quantile plot lie reasonably close to a straight line and show some systematic pattern that is a straight line pattern
Answer #1

First we need to find the mean and s.d of the given data

Mean = 174.4

N = 10

S.d = 39.3226

As the population s.d is unknown here we need to use t distribution

Degrees of freedom is = n-1 = 9

For df 9 and 90% confidence level critical value t from t table is = 1.83

Margin of error (MOE) = t*s.d/√n = 1.93*39.3226/√10 = 22.795

Confidence interval is given by

(Mean - MOE, Mean + MOE)

[151.605, 197.195]